# Reservoir Predictive Path Integral Control for Unknown Nonlinear Dynamics

### v1

@@ Abstract
Neural networks capable of approximating complex nonlinearities have found extensive application in data-driven control of nonlinear dynamical systems. However, fast online identification and control of unknown dynamics remain central challenges. This paper integrates echo-state networks (ESNs)—reservoir computing models implemented with recurrent neural networks—and model predictive path integral (MPPI) control—sampling-based variants of model predictive control—to meet these challenges. The proposed reservoir predictive path integral (RPPI) enables fast learning of nonlinear dynamics with ESN and exploits the learned nonlinearities directly in parallelized MPPI control computation without linearization approximations. The framework is further extended to uncertainty-aware RPPI (URPPI), which leverages ESN uncertainty to balance exploration and exploitation: exploratory inputs dominate during early learning, while exploitative inputs prevail as model confidence grows. Experiments on controlling the Duffing oscillator and four-tank systems demonstrate that URPPI improves control performance, reducing control costs by up to 60% compared to traditional quadratic programming-based model predictive control methods.
Index Terms:
Reservoir Computing, Echo State Network, Model Predictive Control, Path Integral Control, Nonlinear Control, Online Learning
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Achieving high-precision, real-time control of systems with unknown dynamics is challenging in many industrial applications. Standard approaches first estimate the system model through identification, then design model-based control [1, 2]. When the target system is linear, this approach can achieve high control performance based on mature identification and control theory. However, real-world systems exhibit nonlinearities, making linear theory inadequate and degrading performance [3]. To address this, machine learning-based modeling integrated with model-based control has gained much attention [4, 5, 6, 7, 8]. Neural networks (NNs) [4, 5, 6, 7, 8] and Gaussian process regression (GPR) [9, 10] with nonlinear model predictive control (MPC) show promise for nonlinear systems. NNs deliver accurate identification of nonlinear dynamics, whereas GPR enables uncertainty-aware control. This study focuses on integrating data-driven models with MPC to enable real-time control of unknown nonlinear systems.
@@
Although combining machine learning models with nonlinear MPC has shown great promise for controlling nonlinear systems, such approaches often encounter computational challenges in real-time applications. NN-based approaches incur excessive computational overhead in learning model parameters and nonlinear optimization in nonlinear MPC [11]. Similarly, GPR-based methods suffer from computational complexity that scales cubically with the number of training samples [12, 13]. One approach to address these computational challenges is explicit MPC, which pre-computes optimal input mappings [14, 15]. However, required storage and computation demands become prohibitive as the dimensions of the control system’s input and output spaces increase [16]. Consequently, the development of methods that achieve both fast online identification and control remains a fundamental challenge in nonlinear control [17, 18, 15].
@@
Reservoir computing, a framework for efficient neural computation, offers fast training and strong temporal modeling capabilities; among its variants, echo state networks (ESNs) are particularly notable for their simplicity and rapid adaptation [19, 20, 21]. Unlike conventional recurrent NNs, ESNs offer computational efficiency in online training, since only output weights need to be optimized [22]. Using recursive least squares (RLS), the output weights are updated in a computationally efficient manner. Among nonlinear MPC approaches combined with ESN, quadratic programming MPC (QPMPC) has gained attention because of its computational efficiency and real-time capability [23, 24, 25, 26]. This method linearizes a plant model at each control step, casting the optimization problem as a quadratic programming (QP) problem and enabling rapid computation of the optimal control inputs. This approach facilitates online computation of control inputs for nonlinear systems in a computationally efficient manner.
@@
ESN-based QPMPC approaches exhibit two limitations. First, linearizing ESN models and restricting a cost function to a quadratic form both limit the method’s applicability. Because the ESN is linearized at every control step, its nonlinear expressive power cannot be fully exploited, leading to large prediction errors when the plant exhibits strong nonlinearity. Furthermore, constraining the cost function to a quadratic form prevents the inclusion of higher-order terms required by certain control objectives. Thus, developing online control methods that address these limitations remains a pressing research challenge. Second, earlier studies [23, 24, 25, 26] do not discuss the exploration-exploitation trade-off in simultaneous online identification and control. Control inputs that span a wide range of frequencies usually improve identification accuracy, but they can reduce tracking performance, clearly showing this conflict [1]. Thus, practical control methods that automatically balance these two goals in real time constitute a key challenge.
@@
Figure 1: Schematic diagram of the proposed RPPI that integrates ESN and MPPI. (a) Overall architecture of RPPI; (b) ESN model for online identification of nonlinear dynamics, in which only the output weight matrix is trained with RLS; (c) MPPI controller for computing control inputs based on the learned ESN models. In URPPI, the output is predicted using the perturbed output weight matrices so that the controller can compute control inputs that balance exploration and exploitation.
@@
To solve the above challenges, we propose an online control method for nonlinear systems, called reservoir predictive path integral (RPPI), which combines an ESN with model predictive path integral (MPPI) control [27]. MPPI is a sampling-based control method that generates numerous input trajectories and computes their weighted average to determine the control input. This approach directly treats nonlinear dynamics and arbitrary cost functions without requiring explicit optimization solvers [28]. Real-time computation has also been shown to be possible through parallel sampling using GPUs [29]. Furthermore, we propose uncertainty-aware RPPI (URPPI), which incorporates model uncertainty into MPPI input weighting. The model uncertainty is quantified through the precision matrix of the ESN output weight matrix, which is estimated via RLS. This mechanism autonomously balances exploration and exploitation by initially selecting exploratory inputs before model identification converges and later shifting toward exploitative inputs as model confidence increases. The main contributions of this research are summarized as follows:
@@
  • Online MPC for unknown nonlinear dynamics: We propose an RPPI control framework that enables online MPC for strongly nonlinear unknown systems and arbitrary cost functions without linearization.
@@
  • Uncertainty-guided control input: We extend RPPI to a URPPI by injecting the ESN output-weight covariance into the MPPI weighting mechanism, which enables automatic scheduling between exploratory and exploitative control inputs.
@@
  • Experimental validation: We demonstrate the effectiveness of the proposed approach on the Duffing oscillator and four-tank system benchmarks, where URPPI achieves up to 60% lower control cost than ESN-based QPMPC.
The schematic diagram of the proposed method is shown in Fig. 1.
@@
In the following, Section II reviews related work on ESN-based MPC and machine learning approaches to control. Section III presents the problem setting of this research. Next, Section IV introduces the ESN model and its learning algorithm, RLS, and then proposes RPPI and URPPI as methods that integrate ESN with MPPI. In Section V, we conduct numerical simulations to evaluate the performance of the proposed methods. Finally, conclusions and directions for future work are discussed in Section VI.
@@
For a random variable $v$ with density $p_{v}$, the expectation of a function $f$ is denoted by ${\mathbb{E}}_{p_{v}}[f(v)]$. The symbol ${\mathcal{N}}(\mu,\Sigma)$ denotes the multivariate normal distribution with mean $\mu$ and covariance matrix $\Sigma$, and its density function is written as ${\mathcal{N}}(\cdot|\mu,\Sigma)$.
@@ II Related Work
Learning-based control methods face a trade-off between model expressiveness and computational efficiency. NN models, including recurrent NN (RNN), gated recurrent units (GRU), long short-term memory (LSTM), transformer model, input convex NN (ICNN), and graph NN (GNN), offer high expressiveness but require computationally expensive gradient computations for online learning [4, 5, 6, 7, 8, 30, 31, 32]. GPR-based approaches provide uncertainty quantification but suffer from $O(N_{\text{data}}^{3})$ computational complexity, where $N_{\text{data}}$ is the number of training data points, limiting real-time applicability [9, 10]. The ESN architecture balances nonlinear modeling capability with computational efficiency by fixing the input and internal reservoir weights and adapting only the linear readout weights through $O(\hat{N}^{2})$ RLS recursions, where $\hat{N}$ is the reservoir state dimension.
Our control strategy is based on MPPI, a sampling-based framework that enables efficient computation of control inputs for complex and non-differentiable dynamics [27]. We extend this foundation to incorporate model uncertainty. While other uncertainty-aware approaches exist, such as tube-based or risk-aware MPPI [33, 34], our approach utilizes a computationally efficient mechanism. Specifically, we use the covariance matrix of the ESN’s output weights, obtained as a natural byproduct of the RLS learning process, as a measure of parametric uncertainty. As a result, the controller can dynamically adjust the balance between exploration and exploitation.
@@ III Problem Formulation
For such a control system, we consider the problem of finding a control input that minimizes the following cost function $J_{\text{ctrl}}$ at each time $t$:
$\underset{U_{t}\in{\mathcal{U}}}{\text{minimize}}\ J_{\text{ctrl}}(U_{t}),$ (3)
$\begin{aligned} J_{\text{ctrl}}(U_{t})&\coloneqq C\left({\mathbf{y}}_{t},{\mathbf{y}}_{t+1},\ldots{\mathbf{y}}_{t+H}\right)\\ &\quad+\sum_{\tau=t}^{t+H-1}\frac{1}{2}({\mathbf{u}}_{\tau}-{\mathbf{u}}^{\text{ref}})^{\top}R({\mathbf{u}}_{\tau}-{\mathbf{u}}^{\text{ref}}),\end{aligned}$ (4)
where ${\mathcal{U}}\subseteq{\mathbb{R}}^{M\times H}$ is the set of admissible control inputs, and $H\in{\mathbb{N}}$ is a positive constant representing the prediction horizon. Additionally, $R\in{\mathbb{R}}^{M\times M}$ is a positive definite matrix representing input weights, and ${\mathbf{u}}^{\text{ref}}\in{\mathbb{R}}^{M}$ is a vector representing the reference value for the control input. The function $C$ is a cost function related to the time series of the output vectors, expressed using output cost functions $\ell\colon{\mathbb{R}}^{L}\to{\mathbb{R}}$ and $\phi\colon{\mathbb{R}}^{L}\to{\mathbb{R}}$ as follows:
@@
In this paper, we assume that the state vector ${\mathbf{z}}_{t}$ and its dimension $N$, state transition function $\mathbf{F}$}, and output function $\mathbf{G}$} in 2 are unknown, and only the output data ${\mathbf{y}}_{t}$ when input data ${\mathbf{u}}_{t}$ is applied is observed. The goal of this paper is to find a control input $U_{t}$ that minimizes the above cost function while identifying the dynamics of 1 and 2.
@@ IV Proposed Method
   Update the ESN output weight ${\mathbf{W}}^{\mathrm{out}}_{t}$ and the precision matrix ${\mathbf{P}}_{t}$ using RLS (Algorithm 2) with the observed plant output ${\mathbf{y}}_{t}$, reservoir state ${\mathbf{x}}_{t}$, previous output weight ${\mathbf{W}}_{t-1}^{\text{out}}$, and precision matrix ${\mathbf{P}}_{t-1}$.
   Calculate the control input sequence $\hat{U}^{*}_{t}$ with MPPI/UMPPI (Algorithm 3), using previous sequence $\hat{U}_{t}=\hat{U}^{*}_{t-1}$, reservoir state ${\mathbf{x}}_{t}$, output weight ${\mathbf{W}}^{\mathrm{out}}_{t}$ and precision matrix ${\mathbf{P}}_{t}$.
@@
This section presents the proposed framework that integrates ESN with MPPI and uncertainty-aware MPPI (UMPPI), addressing computational and performance limitations in existing approaches. The framework combines ESN’s rapid parameter adaptation capabilities with MPPI’s sampling-based optimization to enable online control of unknown nonlinear systems without linearization approximations. The proposed methods, RPPI and URPPI, operate through the following three-step cycle at each control iteration:
@@
    Update the ESN output weight matrix ${\mathbf{W}}^{\mathrm{out}}_{t}$ and its precision matrix ${\mathbf{P}}_{t}$ using RLS.
@@
An overview of the proposed method is shown in Algorithm 1. The remainder of this section details the components of the proposed framework. Section IV-A describes the ESN architecture and online learning using RLS. Section IV-B explains the MPPI-based control input computation using the ESN model. Section IV-C evaluates the computational complexity of the overall procedure.

### v2

@@ Abstract
Neural networks have found extensive application in data-driven control of nonlinear dynamical systems, yet fast online identification and control of unknown dynamics remain central challenges. To meet these challenges, this paper integrates echo-state networks (ESNs)—reservoir computing models implemented with recurrent neural networks—and model predictive path integral (MPPI) control—sampling-based variants of model predictive control. The proposed reservoir predictive path integral (RPPI) enables fast learning of nonlinear dynamics with ESNs and exploits the learned nonlinearities directly in MPPI control computation without linearization approximations. This framework is further extended to uncertainty-aware RPPI (URPPI), which achieves robust stochastic control by treating ESN output weights as random variables and minimizing an expected cost over their distribution to account for identification errors. Experiments on controlling a Duffing oscillator and a four-tank system demonstrate that URPPI improves control performance, reducing control costs by up to 60% compared to traditional quadratic programming-based model predictive control methods.
@@ I Introduction
Achieving high-precision, real-time control of systems with unknown dynamics is challenging in many industrial applications. Standard approaches first estimate the system model through identification, then design model-based control [1, 2]. When the target system is linear, this approach can achieve high control performance based on mature identification and control theory. However, real-world systems exhibit nonlinearities, making linear theory inadequate and degrading performance [3]. To address this, machine learning-based modeling integrated with model-based control has gained much attention [4, 5, 6, 7, 8]. As such implementations, neural networks (NNs) [4, 5, 6, 7, 8] and Gaussian process regression (GPR) [9, 10] with nonlinear model predictive control (MPC) show promise for nonlinear systems. NNs deliver accurate identification of nonlinear dynamics and GPR enables uncertainty-aware control. This study focuses on integrating data-driven models with MPC to enable online control of unknown nonlinear systems.
@@
While combining machine learning models with nonlinear MPC has shown great promise for controlling unknown nonlinear systems, such approaches often face computational challenges in scenarios demanding fast online learning. In practice, NN-based approaches incur excessive computational overhead in learning model parameters and solving optimization in nonlinear MPC [11]. GPR-based methods typically suffer from computational complexity during identification that scales cubically with the number of training samples [12, 13]. Although explicit MPC mitigates these challenges by pre-computing optimal input mappings, offline strategies are infeasible for unknown systems requiring online model updates [14, 15, 16]. Consequently, the development of methods that achieve both fast online identification and control remains a fundamental challenge in nonlinear control [17, 18, 15].
@@
Reservoir computing, a framework for efficient neural computation, offers fast training and strong temporal modeling capabilities; among its variants, echo state networks (ESNs) are particularly notable for their simplicity and rapid adaptation [19, 20, 21]. Unlike end-to-end trained recurrent NNs, ESNs offer computational efficiency in online training, since the recurrent (reservoir) weights are fixed and only output weights need to be learned [22]. Using recursive least squares (RLS), the output weights are updated with a constant computational cost per step, independent of the number of training samples. Among nonlinear MPC approaches combined with ESN, quadratic programming MPC (QPMPC) has gained attention because of its computational efficiency [23, 24, 25, 26]. This method linearizes a plant model at each control step, casting the optimization problem as a quadratic programming (QP) problem to enable rapid computation of the optimal control inputs. This approach facilitates online computation of control inputs for nonlinear systems in a computationally efficient manner.
@@
However, ESN-based QPMPC approaches exhibit two limitations. First, linearizing ESN models and restricting a cost function to a quadratic form both limit the method’s applicability. Because the ESN is linearized at every control step, its nonlinear expressive power cannot be fully exploited, leading to large prediction errors when the plant exhibits strong nonlinearity. Furthermore, constraining the cost function to a quadratic form prevents the inclusion of higher-order terms required by certain control objectives. Second, earlier studies [23, 24, 25, 26] typically rely on point estimates of model parameters, neglecting the uncertainty inherent in online identification. In online scenarios where data is limited, the estimated model may deviate from the true dynamics. Controllers based solely on such nominal models often degrade control performance when the model is inaccurate. Thus, developing control methods that address these limitations remains a pressing challenge.
@@
Figure 1: Schematic diagram of the proposed RPPI that integrates ESN and MPPI. (a) Overall architecture of RPPI; (b) ESN model for online identification of nonlinear dynamics, in which only the output weight matrix is trained with RLS; (c) MPPI controller for computing control inputs based on the learned ESN models. In URPPI, the output is predicted using the perturbed output weight matrices so that the controller can compute control inputs that minimize the expected cost under model uncertainty.
@@
To address the above challenges, we propose a control method called reservoir predictive path integral (RPPI), which combines an ESN with model predictive path integral (MPPI) control [27]. MPPI is a sampling-based control method that generates numerous input trajectories and computes their weighted average to determine the control input. This approach directly treats nonlinear dynamics and arbitrary cost functions without requiring explicit optimization solvers [28]. Real-time computation has also been shown to be possible through parallel sampling using GPUs [29]. Furthermore, we propose uncertainty-aware RPPI (URPPI), which incorporates model uncertainty into MPPI input weighting. The model uncertainty is quantified using the covariance matrix associated with the RLS estimation of the ESN output weights. This mechanism allows the controller to minimize the expected cost over the distribution of possible system dynamics, leading to control inputs that are robust to identification errors. The schematic diagram of the proposed method is shown in Fig. 1. The main contributions of this research are summarized as follows:
@@
  • Online MPC for unknown nonlinear dynamics: By leveraging ESN-based fast online learning, we propose an RPPI control framework that enables online MPC for strongly nonlinear unknown systems and arbitrary cost functions without linearization.
@@
  • Uncertainty-aware stochastic control: We extend RPPI to a URPPI by injecting the ESN output-weight covariance into the MPPI weighting mechanism, which enables stochastic optimal control that approximately minimizes the expected cost under model uncertainty.
@@
  • Experimental validation: We demonstrate the performance of the proposed approach on the Duffing oscillator and four-tank system benchmarks, where URPPI achieves up to 60% lower control cost than ESN-based QPMPC.
@@
In the following, Section II reviews related work on ESN-based MPC and machine learning-based control. Section III presents the problem setting of this research. Next, Section IV introduces the ESN model and its learning algorithm, RLS, and then proposes RPPI and URPPI as methods that integrate ESN with MPPI. In Section V, we conduct numerical simulations to evaluate the performance of the proposed methods. Finally, conclusions and directions for future work are discussed in Section VI.
@@
For a random variable $v$ following a distribution $p$, the expectation of a function $f$ is denoted by ${\mathbb{E}}_{p}[f(v)]$. The symbol ${\mathcal{N}}(\mu,\Sigma)$ denotes the multivariate normal distribution with mean $\mu$ and covariance matrix $\Sigma$, and its density function is written as ${\mathcal{N}}(\cdot|\mu,\Sigma)$.
@@ II Related Work
Learning-based control methods face a trade-off between model expressiveness and computational efficiency. NN models, including recurrent NN (RNN), gated recurrent units (GRU), long short-term memory (LSTM), transformer model, input convex NN (ICNN), and graph NN (GNN), offer high expressiveness but require computationally expensive gradient computations for online learning [4, 5, 6, 7, 8, 30, 31, 32]. GPR-based approaches provide uncertainty quantification but typically suffer from computational complexity that scales with the accumulated training data points $N_{\text{data}}$ (e.g., $O(N_{\text{data}}^{3})$ for standard GPR), which limits their applicability for long-term online learning [9, 10]. In contrast, the ESN architecture balances nonlinear modeling capability with computational efficiency. By adapting only the linear readout weights via RLS, ESN achieves a constant computational cost of $O(\hat{N}^{2})$ per time step, independent of the number of processed data points, where $\hat{N}$ is the reservoir state dimension.
Reinforcement learning is another powerful approach for controlling systems with unknown dynamics [33, 34]. Probabilistic ensembles with trajectory sampling (PETS) utilizes NN ensembles to minimize expected costs [35], and imitation learning methods approximate stochastic control laws to reduce computational limits [36]. Although these methods demonstrate high performance in episodic settings with abundant offline training data or expert demonstrations, this paper focuses on a different scenario: strictly online control starting from scratch. The proposed URPPI framework leverages ESN’s analytical posterior updates for online stochastic optimal control, enabling online uncertainty quantification and robust adaptation within a single execution.
Our control strategy is based on MPPI, a sampling-based framework that enables efficient computation of control inputs for complex and non-differentiable dynamics [27]. We extend this foundation to incorporate model uncertainty. While other uncertainty-aware approaches exist, such as tube-based or risk-aware MPPI [37, 38], our approach utilizes a computationally efficient mechanism. Specifically, we use the covariance matrix of the ESN’s output weights, obtained as a natural byproduct of the RLS learning process, as a measure of parametric uncertainty. As a result, the controller can consider the distribution of possible models and compute inputs that are robust on average.
@@ III Problem Formulation
For such a control system, we consider the problem of finding a control input sequence $U_{t}\in{\mathbb{R}}^{M\times H}$ that minimizes the following cost function $J_{\text{ctrl}}$ at each time $t$:
$\begin{aligned} J_{\text{ctrl}}(U_{t};{\mathbf{z}}_{t})&\coloneqq C\left({\mathbf{y}}_{t},{\mathbf{y}}_{t+1},\ldots{\mathbf{y}}_{t+H}\right)\\ &\quad+\sum_{\tau=t}^{t+H-1}\frac{1}{2}({\mathbf{u}}_{\tau}-{\mathbf{u}}^{\text{ref}})^{\top}R({\mathbf{u}}_{\tau}-{\mathbf{u}}^{\text{ref}}),\end{aligned}$ (3)
where $H\in{\mathbb{N}}$ is the prediction horizon and the output sequence ${\mathbf{y}}_{t},\ldots,{\mathbf{y}}_{t+H}$ is uniquely determined by the initial state ${\mathbf{z}}_{t}$ and the control input sequence $U_{t}$ via the system dynamics 1 and 2. The input sequence $U_{t}$ consists of control input vectors ${\mathbf{u}}_{\tau}$ for $\tau=t,\ldots,t+H-1$. The matrix $R\in{\mathbb{R}}^{M\times M}$ is a positive definite weight matrix, and ${\mathbf{u}}^{\text{ref}}\in{\mathbb{R}}^{M}$ denotes the reference control input. The function $C$ is a cost function related to the time series of the output vectors, expressed using running cost functions $\ell\colon{\mathbb{R}}^{L}\to{\mathbb{R}}$ and terminal cost function $\phi\colon{\mathbb{R}}^{L}\to{\mathbb{R}}$ as follows:
@@
In this paper, we assume that only the input-output pair $({\mathbf{u}}_{t},{\mathbf{y}}_{t})$ is observable at each time step $t$, while the state transition function $\mathbf{F}$} and output function $\mathbf{G}$} in 1 and 2 are unknown. To account for the lack of knowledge, we formulate the control problem as a stochastic MPC problem. We address the problem of finding a control input sequence $U_{t}$ that minimizes the expected cost with respect to the parameter ${\mathbf{w}}_{t}$ representing the model uncertainty:
$\underset{U_{t}\in{\mathcal{U}}}{\text{minimize}}\ {\mathbb{E}}_{\mathcal{P}_{t}}[J_{\text{ctrl}}(U_{t})],$ (5)
$\text{subject to}\ {\mathbf{x}}_{t+1}=\hat{{\mathbf{F}}}({\mathbf{x}}_{t},{\mathbf{u}}_{t};{\mathbf{w}}_{t}),$ (6)
[…]
Here, the system dynamics are approximated by a surrogate model $\hat{\mathbf{F}}$ and $\hat{\mathbf{G}}$ parameterized by ${\mathbf{w}}_{t}$. The symbol $\mathcal{P}_{t}$ denotes the probability distribution of the parameter ${\mathbf{w}}_{t}$. The variable ${\mathbf{x}}_{t}$ denotes the state of the surrogate model, and $\hat{{\mathbf{y}}}_{t}\in{\mathbb{R}}^{L}$ denotes its output. The cost function $J_{\text{ctrl}}(U_{t})$ in 5 is evaluated along the trajectory generated by this surrogate model. The symbol ${\mathcal{U}}\subseteq{\mathbb{R}}^{M\times H}$ is the set of admissible control inputs. As reviewed in Ref. [39], this stochastic formulation enables the controller to account for model uncertainty, leading to a robust control strategy that reduces the risk of performance degradation due to identification errors. The goal of this paper is to solve this stochastic MPC problem online while simultaneously identifying the system dynamics from observed data.
Solving 5 presents two major challenges. First, obtaining the surrogate model of the system dynamics, including the uncertainty distribution $\mathcal{P}_{t}$, from limited data online is difficult for general nonlinear models. Second, minimizing the cost function is computationally intractable, as it results in a non-convex optimization problem when the model is nonlinear and the cost function includes high-order terms. Our aim is to overcome these challenges.
@@ IV Proposed Method
   Update the ESN output weight ${\mathbf{W}}^{\text{out}}_{t}$ and the precision matrix ${\mathbf{P}}_{t}$ using RLS (Algorithm 2) with the observed plant output ${\mathbf{y}}_{t}$, reservoir state ${\mathbf{x}}_{t}$, previous output weight ${\mathbf{W}}_{t-1}^{\text{out}}$, and precision matrix ${\mathbf{P}}_{t-1}$.
   Calculate the control input sequence $\hat{U}^{*}_{t}$ with MPPI/UMPPI (Algorithm 3), using previous sequence $\hat{U}_{t}=\hat{U}^{*}_{t-1}$, reservoir state ${\mathbf{x}}_{t}$, output weight ${\mathbf{W}}^{\text{out}}_{t}$ and precision matrix ${\mathbf{P}}_{t}$.
@@
This section presents a framework that integrates the ESN with MPPI. To overcome the difficulty of online probabilistic modeling, we employ an ESN trained with RLS. Distinct from standard deterministic approaches, we utilize the precision matrix estimated by RLS to quantify the posterior variance of the model parameters, providing a lightweight and analytical expression for model uncertainty. To address the intractability of non-convex optimization, we adopt the MPPI framework. This method approximates the optimal control problem as a free-energy minimization that can be solved efficiently via importance sampling without linearization approximations. The integration of these components results in the URPPI algorithm, providing a unified framework for online stochastic optimal control of unknown nonlinear systems.
The proposed methods, RPPI and URPPI, operate through the following three-step cycle at each control iteration:
@@
    Update the ESN output weight matrix ${\mathbf{W}}^{\text{out}}_{t}$ and its precision matrix ${\mathbf{P}}_{t}$ using RLS.
@@
An overview is shown in Algorithm 1. The remainder of this section details the components of the proposed framework. Section IV-A describes the ESN architecture and online learning using RLS. Section IV-B explains the MPPI-based control using the ESN model. Section IV-C evaluates the computational complexity of the overall procedure.